Activity No. 1: Multicab Ride!
Going to a far barangay in Gingoog City means riding a multicab. The standard fare in riding a multicab is Php 7 for the first 5 km as a flag down rate plus Php 1 peso for every 1 kilometer. How much is Nathan going to pay if he goes to Medina?
Complete the table:
5 km - Lunao 8 km - agay 11 km - Minlagas 13 km - San Luis 18 km - Medina 23 km

Directions: Consider the situation below and answer the question that follows. 5pts
1. What is the dependent variable?
2. What is the independent variable?
3. Based on the completed table, would the relation represent a line?
4. What is the y-intercept?
5. What is the slope?

Answer by ewatrrr(24785)

 
Hi,
P(x,y): (5,7), (8,10), (11,13), (23,25)  x = mi and y = cost

1. What is the dependent variable? y (the cost)

2. What is the independent variable? x  (the mi)

3. Based on the completed table, 
would the relation represent a line? Yes   y = x + 2

4. What is the y-intercept? 2

5. What is the slope? 1

  y = (x-5) + 7    0r y = x+2

Wish You the Best in your Studies.